OpenStudy (anonymous):

Write an equation of the line that passes through the point (–4, –1) and isperpendicular to the line 2x + 7y = 14.

OpenStudy (anonymous):

First isolate y and find the slope of the equation 2x + 7y = 14

OpenStudy (anonymous):

That is perpendicular *

OpenStudy (anonymous):

Do you want the answer or know how to do it?

OpenStudy (anonymous):

answer please .

OpenStudy (anonymous):

y = -2/7 x + b -1 = 8/7 + b b = -15/7 equation is y = -2/7 x - 15/7

OpenStudy (anonymous):

Thanks moneybird . ^_^